# Inverse-Free Fast Natural Gradient Descent Method for Deep Learning

### v1

@@ IV-C Per-sample Gradient
where $\mathcal{Z}$ is the gradient of the output patches $\mathcal{Y}$. In distributed learning, we can only transmit the input patches and the gradient of output to reduce the communication burden. Due to the summation operator, we can’t employ the identity equation to reconstruct $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ as Eq. 12. Nonetheless, as the number of parameters in convolutional layers is much smaller than that of fully connected layers, the computation of $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ is typically affordable.
@@ IV-D Setting of Damping
V Experiments
@@
In this section, we compare FNGD with prevailing first-order methods, such as SGD and AdamW [12], as well as second-order methods, like KFAC, shampoo, and Eva. We examine their performance on the following two tasks: image classification and machine translation. Each algorithm was executed with the optimal hyperparameters determined through a grid search. For image classification, we decay the learning rate by a factor of 0.1 at 50% and 75% of the training epochs. For machine translation, we keep the learning rate constant. In the case of KFAC and Shampoo, we set the frequency for updating second-order statistics to $T_{1}=10$ and the frequency for inverting to $T_{2}=100$. For Eva, we update the second-order statistics during every iteration.
@@
We only utilize second-order statistics to precondition the gradient of convolutional layers and fully connected layers. For BatchNorm layers, LayerNorm layers, and embedding layers, we directly use the gradient descent direction. When implementing the KFAC algorithm, we follow the suggestion in [18, 6] that employs eigenvalue decomposition on Kronecker factors to compute the inverse, which has been shown to yield higher test accuracy compared to directly inverting. For the implementation of Shampoo and Eva, we use the publicly available code [^2] https://github.com/lzhangbv/eva. Our experiments were run on GeForce RTX 3060Ti GPUs using Pytorch.
(a) ResNet-32
(b) ResNet-110
@@
Figure 5: The convergence curves of FNGD, SGD-m, KFAC, Shampoo, and Eva on ResNet-32 and ResNet-110 with the CIFAR-10 dataset.
@@
Figure 6: The convergence curves of FNGD, SGD-m, KFAC, Shampoo, and Eva on ResNet-18 and ResNet-34 with the CIFAR-100 dataset.
V-A Image Classification
We first evaluate the effectiveness and time efficiency of our method on image classification tasks. In order to examine our method on networks with different widths and depths, we run experiments on four ResNet [9] models: ResNet-32 and ResNet-110 with the CIFAR-10 dataset, and ResNet-18 and ResNet-34 with the CIFAR-100 dataset. The two datasets both have 50,000 training samples and 10,000 test samples. In our experiments, the first-order method SGD with momentum 0.9 (SGD-m) was run for 200 epochs, while second-order methods were run for 100 epochs. We set a batch size of 128 for all algorithms.
We present the convergence curves of FNGD and the other mentioned algorithms for CIFAR-10 and CIFAR-100 in Fig. 5 and Fig. 6, respectively. One can see that, for the four image classification tasks, FNGD can achieve comparable convergence and generalization performance when compared to other second-order methods. In comparison to the first-order method, FNGD can achieve convergence in only half the number of iterations. Specifically, we take the ResNet-110 on CIFAR-10 as an example to present some detailed statistical results in Tab. II. To achieve the target accuracy of 93.5%, FNGD requires the fewest iterations, nearly half that needed by SGD-m. Furthermore, FNGD can achieve the target accuracy within the shortest running time. Compared to SGD-m, KFAC increases the running time by 18.3%, while FNGD reduces it by 17.5%. The results of Eva are not listed in Tab. II as its maximum test accuracy is 93.45%, which doesn’t reach the target.
TABLE II: Comparison of FNGD and other methods in terms of running time when reaching the target accuracy of 93.5% for ResNet-110 on CIFAR-10.
Method SGD-m KFAC Shampoo FNGD
Epoch 151 80 78 77
Time (s) 4299 5086 4272 3549
Time Gap 0% +18.3% -0.6% -17.5%
@@ V-A Image Classification
CIFAR-10 ResNet-32 1$\times$ 1.71$\times$ 1.51$\times$ 1.83$\times$ 1.23$\times$
ResNet-110 1$\times$ 2.24$\times$ 1.92$\times$ 2.35$\times$ 1.58$\times$
CIFAR-100 ResNet-18 1$\times$ 2.79$\times$ 1.62$\times$ 1.75$\times$ 1.36$\times$
ResNet-34 1$\times$ 2.76$\times$ 1.56$\times$ 1.56$\times$ 1.42$\times$
In Tab. III, we compare the per-epoch training time of FNGD and other algorithms. It is evident that FNGD exhibits the shortest per-epoch training time among all the evaluated second-order methods. On average, the per-epoch training time of FNGD is 1.37$\times$ longer than that of SGD. When compared to KFAC, Shampoo, and Eva, FNGD can achieve speedup factors of up to 2.05$\times$, 1.22$\times$, and 1.49$\times$, respectively. Note that the relative time cost of Eva is higher than what is reported in [25]. This is because the batch size we utilize is much smaller than the 1024 mentioned in [25], which results in more iterations per epoch. Consequently, there will be more statistical information computations and preconditioning operators.
V-B Machine Translation
@@
For natural language processing tasks, SGD performs much worse than AdamW as demonstrated in [23]. Therefore, we conducted comparative experiments with AdamW. We didn’t include Eva in our experiments as its effectiveness for the Transformer has not been confirmed in [25]. We run all the algorithms for 100 epochs with a batch size of 64.
Our results are shown in Fig. 7. It is demonstrated that FNGD yields the highest BLEU score on the test dataset. Specifically, the BLEU score of FNGD is 24 higher than that of AdamW, 11 higher than that of Shampoo, and 4 higher than that of KFAC. Furthermore, FNGD outperforms these second-order methods in terms of end-to-end training time. FNGD can achieve a comparative time cost compared to AdamW. In comparison to KFAC, FNGD is approximately 2.4$\times$ faster. Moreover, there is a large time gap between Shampoo and FNGD, which differs from the situation with the ResNet models. This is because the dimensions of layers in the Transformer are no less than 512, thereby significantly increasing the computational complexity associated with inverting.
@@
V-C Time Analysis
In order to have a thorough understanding of the time cost of FNGD, we provide a detailed analysis of the time cost of each step within FNGD in Fig. 8. We present the analysis on two different types of networks, i.e. , ResNet and Transformer. As in SGD-m, the training process involves three primary steps: forward pass, backward pass, and parameter update. The FNGD and AdamW have an additional preconditioning step. For the ResNet model, we have to extract the input feature of convolutional layers into patches to compute the per-sample gradients.
In Fig. 8, we can see that the backward time in FNGD is less than the standard time cost in SGD-m. This results from our strategy of efficiently computing per-sample gradients. As mentioned in Sec. IV-C, we propose to make use of Autograd to compute the gradient of modules’ output. It can be seen as a substep in the SGD-m process of computing the gradient of modules’ parameters.
The patch extraction operator makes up 24% of the total time cost of FNGD as shown in LABEL:fig:sub_resnet110. It matches the time required for the preconditioning step, which involves computing the per-sample gradients and performing the weighted sum. The computational complexity of the extraction operator is dependent on the size of the feature map. Consequently, for deep networks with plenty of wide convolutional layers, the extraction operator becomes time-consuming. This analysis can provide an explanation of the results in Tab. III. When compared to SGD-m, the relative time costs of FNGD on ResNet-110 and ResNet-34 are a little higher than those on ResNet-32 and ResNet-18. This can be attributed to their wider and deeper network structure.
In the Transformer architecture, there is no need for the time-consuming patch extraction operator due to the absence of convolutional layers. As a result, the time cost of FNGD closely approaches that of AdamW.
(a) ResNet-110
@@ V-C Time Analysis
Figure 8: Analysis of the time cost of FNGD and first-order methods on two types of model structures.
VI Conclusion

### v2

@@ IV-C Per-sample Gradient
where $\mathcal{Z}$ is the gradient of the output patches $\mathcal{Y}$. Due to the summation operator, we can’t employ the identity equation to reconstruct $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ as Eq. 12. Nonetheless, as the number of parameters in convolutional layers is much smaller than that of fully connected layers, the computation of $\mathbf{U}_{l}^{\text{T}}\mathbf{U}_{l}$ is typically affordable.
After deriving these coefficients, to mitigate the significant memory overhead associated with storing $\mathbf{U}_{l}$, we propose to apply weighted coefficients directly on inputs, and then we can derive the preconditioned gradient through the gradient computation tools in PyTorch, by using weighted inputs and output gradients. This approach also cancels out the time-consuming patch-extracting operator. Therefore, the additional computational complexity introduced by FNGD over SGD is primarily due to the input weighting step with $\operatorname{O}(MD)$ complexity as listed in Tab. I.
@@ IV-D Setting of Damping
V Convergence
We follow the framework established in [29] to provide a theoretical convergence analysis. For simplicity, we consider the single-output case with squared error loss, while the multiple-output case can be extended similarly. Furthermore, we focus on the full-batch case.
Given a dataset $\{\mathbf{x}_{m},y_{m}\}_{i=1}^{M}$ with $\mathbf{x}_{m}\in\mathbb{R}^{d}$ and $y_{m}\in\mathbb{R}$, we have the network output vector denoted as $\mathbf{v}(\mathbf{w})=[\operatorname{f}(\mathbf{w},\mathbf{x}_{1}),\cdots,\operatorname{f}(\mathbf{w},\mathbf{x}_{M}))]^{\text{T}}$, and the squared error loss can be denoted as:
$\min_{\mathbf{w}}\frac{1}{2}\|\mathbf{v}(\mathbf{w})-\mathbf{y}\|^{2}_{2},$
where $\mathbf{y}=[y_{1},y_{2},\cdots,y_{M}]^{\text{T}}$. We denote the Jacobian matrix as $\mathbf{J}(\mathbf{w})=[\frac{\partial\operatorname{f}(\mathbf{w},\mathbf{x}_{1})}{\partial\mathbf{w}},\cdots,\frac{\partial\operatorname{f}(\mathbf{w},\mathbf{x}_{M})}{\partial\mathbf{w}}]\in\mathbb{R}^{N\times M}$. Consequently, the averaged gradient vector $\mathbf{g}(\mathbf{w})\in\mathbb{R}^{N}$ is given by $\frac{1}{M}\mathbf{J}(\mathbf{w})(\mathbf{v}(\mathbf{w})-\mathbf{y})$.
In FNGD, the fisher information matrix is block-diagonal approximated, i.e. $\mathbf{F}=\frac{1}{M}\tilde{\mathbf{J}}\tilde{\mathbf{J}}^{\text{T}}$, where $\tilde{\mathbf{J}}\in\mathbb{R}^{N\times ML}$ is a variant of $\mathbf{J}$. This variant is defined as $\mathbf{J}=\tilde{\mathbf{J}}\mathbf{K}$, with $\mathbf{K}=[\mathbf{I}_{M},\mathbf{I}_{M},\cdots,\mathbf{I}_{M}]^{\text{T}}\in\mathbb{R}^{ML\times M}$. Here, $L$ denotes the number of network layers.
We make two assumptions that suffice to prove the global convergence of FNGD. The first assumption is in line with the coefficient-sharing technique in FNGD. The second assumption requires that $\mathbf{J}$ is stable enough for small perturbations around the initialization, which guarantees that the network is close to a linearized network.
Assumption 1.
The Gram matrix $\mathbf{G}=\tilde{\mathbf{J}}^{\text{T}}\tilde{\mathbf{J}}$ is constant across iterations, and $\mathbf{G}$ is a positive definite matrix, with $0<\lambda_{\min}\leq\operatorname{\lambda}(\mathbf{G})\leq\lambda_{\max}<+\infty$.
Assumption 2.
For all parameters $\mathbf{w}$ that satisfy $\|\mathbf{w}-\mathbf{w}_{0}\|_{2}\leq\frac{\sqrt{\lambda_{\max}L}}{\lambda_{\min}}\|\mathbf{y}-\mathbf{v}(\mathbf{w}_{0})\|_{2}$, we have $\|\mathbf{J}(\mathbf{w})-\mathbf{J}(\mathbf{w}_{0})\|_{2}\leq\frac{\lambda_{\min}}{2\sqrt{\lambda_{\max}}}$.
Theorem 1.
Let Assumptions 1 and 2 hold. Suppose we optimize with FNGD using a damping value $\lambda=\frac{\lambda_{\min}}{M}$ and a small enough learning rate $\eta\leq\tilde{\eta}$, we have $\|\mathbf{v}_{k}-y\|_{2}^{2}\leq(1-\eta)^{k}\|\mathbf{v}_{0}-y\|_{2}^{2}$.
The detailed proof can be found in the Appendix. Although the optimization landscape is non-convex, FNGD with a constant learning rate and damping value enjoys at least a linear convergence rate. With a larger minimum eigenvalue $\lambda_{min}$, it is feasible to employ a larger learning rate so that FNGD can converge within fewer iterations.
VI Experiments
@@
In this section, we compare FNGD with prevailing first-order methods, such as SGD and AdamW [15], as well as second-order methods, like KFAC, shampoo, and Eva. We examine their performance on the following two tasks: image classification and machine translation. Each algorithm was executed with the optimal hyperparameters determined through a grid search. In the case of KFAC and Shampoo, we set the frequency for updating second-order statistics to $T_{1}=10$ and the frequency for inverting to $T_{2}=100$. For Eva, we follow the setting described in [30] to update the second-order statistics at each iteration.
@@
For all the algorithms mentioned above, we only utilize second-order statistics to precondition the gradient of convolutional layers and fully connected layers. For other trainable layers including BatchNorm layers, LayerNorm layers, and Embedding layers, we directly adopt the gradient descent direction. When implementing the KFAC algorithm, we follow the suggestion in [21, 8] that employs eigenvalue decomposition on Kronecker factors to compute the inverse, which has been shown to yield higher test accuracy compared to directly inverting. For the implementation of Shampoo and Eva, we use the publicly available code [^2] https://github.com/lzhangbv/eva. Our experiments were run on GeForce RTX 3080 GPUs using PyTorch.
VI-A Image Classification
We first evaluate our method’s effectiveness and time efficiency on image classification tasks. To evaluate our method across Convolutional Neural Networks (CNNs) with varying widths and depths, we run experiments using five models: ResNet-32 and VGG-11 on CIFAR-10 dataset, ResNet-18 and ResNet-34 on CIFAR-100 dataset, and ResNet-50 on large-scale ImageNet dataset. In our experiments of CIFAR-10 and CIFAR-100, the first-order method SGD with momentum 0.9 (SGD-m) was run for 200 epochs, while second-order methods were run for 100 epochs. We decay the learning rate by a factor of 0.1 at 50% and 75% of the training epochs. Regarding ImageNet, the number of epochs was set to 90 for SGD-m and 55 for second-order methods. We set a batch size of 128 for all algorithms.
We present the optimization curves of FNGD and the other mentioned algorithms for CIFAR-10 and CIFAR-100 in Fig. 5 and Fig. 6, respectively. One can see that, for the four image classification tasks, FNGD can achieve comparable convergence and generalization performance when compared to other second-order methods. In terms of time efficiency, FNGD outperforms the other algorithms. It is noticeable that in the majority of cases, KFAC tends to be slower than SGD-m. In contrast, FNGD achieves comparable or superior accuracy to SGD-m in merely half the time.
This time efficiency contributes to the efficient training of FNGD. In Tab. II, we take the ResNet-32 on CIFAR-10 as an example to demonstrate the optimization advantage of FNGD. As indicated in Tab. II, FNGD achieves a 32.6% reduction in training time compared to SGD-m to attain the target accuracy of 92.5%. In contrast, Eva and Shampoo exhibit increase in training time of 17.3% and 1.7%, respectively.
(a) VGG-11
(b) ResNet-32
@@ VI-A Image Classification
Figure 5: The optimization curves of FNGD, SGD-m, KFAC, Shampoo, and Eva on ResNet-32 and VGG-11 with the CIFAR-10 dataset.
@@
Figure 6: The optimization curves of FNGD, SGD-m, KFAC, Shampoo, and Eva on ResNet-18 and ResNet-34 with the CIFAR-100 dataset.
TABLE II: Comparison of FNGD and other methods in terms of running time when reaching the target accuracy of 92.5% for ResNet-32 on CIFAR-10.
Method SGD-m KFAC Shampoo Eva FNGD
Epoch 112 75 77 76 75
Time (s) 1677 1542 1706 1967 1131
Time Gap 0% -8.1% +1.7% +17.3% -32.6%
@@
CIFAR-10 VGG-11 1$\times$ 2.73$\times$ 2.09$\times$ 1.75$\times$ 1.03$\times$
ResNet-32 1$\times$ 1.52$\times$ 1.57$\times$ 1.84$\times$ 1.06$\times$
CIFAR-100 ResNet-18 1$\times$ 2.18$\times$ 1.48$\times$ 1.40$\times$ 1.03$\times$
ResNet-34 1$\times$ 2.22$\times$ 1.69$\times$ 1.26$\times$ 1.05$\times$
We give a detailed time efficiency comparison of FNGD and other algorithms in Tab. III. It is evident that FNGD exhibits the shortest per-epoch training time among all the evaluated second-order methods, nearly matching the performance of SGD-m. On average, the per-epoch training time of FNGD is 1.04$\times$ longer than that of SGD. When compared to KFAC, Shampoo, and Eva, FNGD can achieve speedup factors of up to 2.07$\times$, 1.64$\times$, and 1.50$\times$, respectively. Note that the relative time cost of Eva is higher than what is reported in [30]. This is because the batch size we utilize is much smaller than the 1024 mentioned in [30], which results in more iterations per epoch. Consequently, there will be more statistical information computations and preconditioning operators.
For the larger-scale dataset ImageNet, we evaluate FNGD’s performance on ResNet50 as shown in Fig. 7. It is clear that the training time of FNGD is nearly half of that of SGD, and FNGD is the fastest algorithm to achieve the target accuracy of 75%. For the distributed implementation of FNGD, we can first perform the weighted sum of per-sample gradients on each server. Subsequently, these distributed preconditioned gradients will be aggregated. In this way, there are no additional communication costs introduced by FNGD.
Figure 7: Generalization performance for ResNet-50 on ImageNet.
VI-B Machine Translation
@@
For natural language processing tasks, SGD performs much worse than AdamW as demonstrated in [26]. Therefore, we included comparative experiments with AdamW. We didn’t include Eva in our experiments as its effectiveness for the Transformer has not been confirmed in [30]. We run all the algorithms for 100 epochs with a batch size of 64.
Our results are shown in Fig. 8. It is demonstrated that FNGD yields the highest BLEU score on the test dataset. Specifically, the BLEU score achieved by FNGD surpasses that of SGD-m by nearly 35 points, and FNGD outperforms AdamW by 24 points, Shampoo by 11 points, and KFAC by 4 points. Furthermore, FNGD outperforms these second-order methods in terms of end-to-end training time. FNGD can achieve a similar time cost as both SGD-m and AdamW. In comparison to KFAC, FNGD is approximately 2.4$\times$ faster. Moreover, there is a remarkable time gap between Shampoo and FNGD, which differs from the situation with CNN models. This is attributed to the fact that the dimensions of layers in the Transformer are no less than 512, which significantly increases the computational complexity associated with inverting.
@@
VI-C Time Analysis
In order to have a thorough understanding of the time cost of FNGD, we provide a detailed analysis of the time cost of each step within FNGD in Fig. 9. We present the analysis on two different types of networks, i.e. , ResNet and Transformer. As in SGD-m, the training process involves three primary steps: forward pass, backward pass, and parameter update. The FNGD and AdamW have an additional preconditioning step.
In Fig. 9, we can see that the backward time in FNGD is less than the standard time cost in SGD-m. This results from our strategy of efficiently computing per-sample gradients. As mentioned in Sec. IV-C, we propose to make use of Autograd to compute the gradient of modules’ output. It can be seen as a substep of the SGD-m backward process.
The total time spent on the backward and preconditioning processes of FNGD is comparable to the backward time of SGD-m, resulting in the overall computational complexity of FNGD approaches that of SGD-m. For the wider Transformer structure, its total time slightly exceeds that of the standard backward, as the computational cost of the preconditioning step is related to the layer dimension.
(a) ResNet-32
@@ VI-C Time Analysis
Figure 9: Time Analysis of FNGD and first-order methods on two types of model structures.
VI-D Ablation Study
We perform an ablation study to validate the necessity of the three key components in FNGD: coefficient-sharing, efficient precondition for acceleration, and damping determination. Without employing damping determination, we assigned a constant damping value of 0.3 to all layers in this experiment. Without applying efficient precondition, we start by performing the standard backward pass, leveraging hooks to derive the per-sample gradients, and subsequently compute the matrix-vector multiplication. The results are demonstrated in Tab. IV.
It is demonstrated that any of the three key components is crucial for FNGD. Specifically, coefficient-sharing can reduce the training time by half with a small increase in accuracy. By employing Autograd on module output and performing efficient weighed-sum, a slight speedup can be achieved compared with using hooks. Without the damping determination strategy described in Sec. IV-D, there will be an accuracy degeneration, particularly in the case of ResNet-18.
TABLE IV: Ablation study on FNGD. We train ResNet-110 on Cifar-10, and ResNet-18 on CIFAR-100. The first row represents test accuracy, and the second row represents the relative per-epoch training time.
Network FNGD w/o sharing w/o acceleration w/o damping
ResNet-32 93.10% 92.75% - 92.63%
1.0$\times$ 2.0$\times$ 1.3$\times$ -
ResNet-18 77.37% 77.06% - 75.51%
1.0$\times$ 2.1$\times$ 1.2$\times$ -
VI-E Hyper-parameter
We explore the impact of different hyperparameter settings on the generalization performance of FNGD in comparison with KFAC. Our particular focus lies on the damping value. With the damping determination strategy, the damping value tuned in FNGD is referred to as the $\alpha$ in Eq. 14. We perform the experiment by training ResNet-18 on CIFAR-100. For fairness, the other hyperparameters are fixed for both FNGD and KFAC.
We represent the accuracy of FNGD and KFAC with different damping values in Tab. V. Specifically, by varying the damping value from 0.01 to 0.001, the mean accuracy of FNGD reaches 77.22%, while KFAC exhibits a lower mean value of 77.00%. Moreover, the highest accuracy 77.56% of FNGD is much higher than that of KFAC, 77.22%. In general, FNGD is less likely to require tedious fine-tuning to achieve high accuracy.
TABLE V: Hyper-parameter study on damping value by training ResNet-18 on CIFAR-100.
Method 0.01 0.008 0.005 0.003 0.001
FNGD 77.24% 77.09% 77.56% 77.34% 76.86%
KFAC 77.22% 76.96% 77.00% 76.71% 77.10%
VII Conclusion